A 250.0-mg sample of an organic weak acid is dissolved in an appropriate solvent and titrated with 0.0556 M NaOH, requiring 32.58 mL to reach the end point. Determine the compound’s equivalent weight.


1
Expert's answer
2021-06-10T03:32:28-0400

"\star" Moles of titrant

The moles of NaOH needed to reach the end point are:

"M_b\times V_b= 0.0556M\times 0.03258L"

"=0.00181" moles of NaOH


"\star" Wt. of analyte=250 X10-3 g


This gives an equivalent weight of analyte:


E .W. = (g analyte/mole titrant)


="250\times 10^{-3} \over 0.00181"

="138.1" g/mol